OpenStudy (anonymous):

please differentiate this: y=x/the square root of x+1 ..

mathslover (mathslover):

\[\large{\frac{d}{dx}(\frac{x}{\sqrt{x+1}})}\]

OpenStudy (anonymous):

use quotient rule, then chain rule

OpenStudy (anonymous):

hmm .. im a begginer only .. i dont know how to use that ,,

mathslover (mathslover):